### v2

# Low-Rank Orthogonalization for Large-Scale Matrix Optimi- zation with Applications to Foundation Model Training

Abstract
Neural network (NN) training is inherently a large-scale matrix optimization problem, yet the matrix structure of NN parameters has long been overlooked. Recently, the optimizer Muon (Jordan et al., ), which explicitly exploits this structure, has gained significant attention for its strong performance in foundation model training. A key component contributing to Muon’s success is matrix orthogonalization. In this paper, we propose low-rank orthogonalization, which performs orthogonalization by leveraging the low-rank nature of gradients during NN training. Building on this, we introduce low-rank matrix-signed gradient descent (MSGD) and a low-rank variant of Muon. Numerical experiments demonstrate the superior performance of low-rank orthogonalization, with low-rank Muon achieving promising results in GPT-2 and LLaMA pretraining—surpassing the carefully tuned vanilla Muon on tasks with large model sizes. Theoretically, we establish the iteration complexity of low-rank MSGD for finding an approximate stationary solution, and the iteration complexity of low-rank Muon for finding an approximate stochastic stationary solution under heavy-tailed noise. The code to reproduce our numerical experiments is available at https://github.com/dengzhanwang/Low-rank-Muon.
Keywords: Orthogonalization, Muon, foundation model training, iteration complexity, heavy-tailed noise
1 Introduction
Training neural networks (NNs) (LeCun et al., 2015), particularly recent foundation models, has consistently posed challenging large-scale optimization problems. Over the past decade, NN training has been dominated by vector-variate optimization methods—including SGD (Bottou, 2010), AdaGrad (Duchi et al., 2011), RMSprop (Hinton et al., 2012), Adadelta (Zeiler, 2012), Adam (Kingma and Ba, 2015), and AdamW (Loshchilov and Hutter, 2019). Nonetheless, these methods disregard the inherent matrix structure of NN parameters—such as those in multi-layer perceptrons (Rosenblatt, 1958), convolutional layers (LeCun et al., 1989), and the query, key, and value projections in attention mechanisms (Vaswani et al., 2017).
Recently, a shift has occurred: optimization methods that exploit matrix structure are receiving increasing attention and have begun to demonstrate strong performance in foundation model training (Jordan et al., ; Pethick et al., 2025). These methods focus on solving the matrix optimization problem:
$\min_{X\in\mathbb{R}^{m\times n}}f(X).$ (1)
In particular, Shampoo, as developed in (Anil et al., 2020; Gupta et al., 2018), applies left and right preconditioning matrices and updates them as follows:
$X^{k+1}=X^{k}-\eta_{k}(L^{k})^{-1/4}G^{k}(R^{k})^{-1/4},$ (2)
where $\eta_{k}>0$ is the step size, $G^{k}\in\mathbb{R}^{m\times n}$ denotes the (stochastic) gradient of $f$ at $X^{k}$, and $L^{k}\in\mathbb{R}^{m\times m}$ and $R^{k}\in\mathbb{R}^{n\times n}$ are the left and right preconditioning matrices, respectively. Shampoo updates the left and right preconditioners $\{L^{k}\}$ and $\{R^{k}\}$ using the second-order statistics of the accumulated gradients, similarly to AdaGrad, and has shown comparable performance to popular vector-variate optimizers such as Adam and AdamW on foundation model training. Following Shampoo, other matrix-variate optimizers that use two-sided preconditioners, such as CASPR (Duvvuri et al., 2024) and SOAP (Vyas et al., 2025), were also developed. A preconditioned Riemannian gradient descent method was developed in (Bian et al., 2024) for low-rank matrix recovery, which adopts only the diagonal part of the Shampoo preconditioners. Moreover, one-sided preconditioned variants of Shampoo were developed in (An et al., 2025; Xie et al., 2025).
In addition to Shampoo and its variants, another matrix-variate optimizer, Muon (Jordan et al., ), has attracted significant attention for outperforming standard optimizers such as Adam and AdamW in foundation model training (AI et al., 2025). At each iteration, Muon performs the update:
$M^{k}=(1-\theta_{k-1})M^{k-1}+\theta_{k-1}G(X^{k};\xi^{k}),\quad X^{k+1}=X^{k}-\eta_{k}\mathrm{msgn}(M^{k}),$ (3)
where $G(\cdot;\xi)$ denotes the stochastic gradient of $f(\cdot)$, and $\mathrm{msgn}(M^{k})=U^{k}(V^{k})^{T}$ denotes the matrix sign of $M^{k}$, with $U^{k}$ and $V^{k}$ containing the left and right singular vectors of $M^{k}$, respectively. The matrix sign computation is often referred to as matrix orthogonalization, because calculating $\mathrm{msgn}(M)$ is equivalent to finding the (semi-)orthogonal matrix closest to $M$ with respect to the Frobenius norm (see, e.g., (Bernstein and Newhouse, 2024, Proposition 4)). Muon’s empirical success has sparked significant research interest, including efforts to understand its relationship with other algorithms, establish its convergence guarantees, and propose new variants (see, e.g., (An et al., 2025; Cesista, 2025; Chen et al., 2025; Glentis et al., 2025; Kovalev, 2025; Lau et al., 2025; Li and Hong, 2025; Liu et al., 2025; Ma et al., 2024; Pethick et al., 2025; Riabinin et al., 2025; Sato et al., 2025; Sfyraki and Wang, 2025; Shen et al., 2025)). A popular interpretation of Muon is from the perspective of a linear minimization oracle with respect to the spectral norm (e.g., see (Bernstein and Newhouse, 2024; Cesista, 2025; Glentis et al., 2025; Kovalev, 2025; Lau et al., 2025; Riabinin et al., 2025)). That is, the matrix sign computation in (3) can be recast as $\mathrm{msgn}(M^{k})={\arg\min}_{\|\Delta\|\leq 1}\{\langle M^{k},\Delta\rangle\}$, where $\|\cdot\|$ denotes spectral norm. Based on this interpretation, algorithmic designs leveraging general matrix-induced norms $\|\cdot\|_{p\to q}$ have been discussed in (Bernstein and Newhouse, 2024; Cesista, 2025; Glentis et al., 2025; Riabinin et al., 2025). In addition, Muon is also connected to earlier algorithms and can be viewed as a special case of Shampoo, despite not explicitly using preconditioners. As discussed in (Jordan et al., ), by taking $L^{k}=G^{k}(G^{k})^{T}$ and $R^{k}=(G^{k})^{T}G^{k}$ in (2), the Shampoo updates reduce to the matrix-signed update: $X^{k+1}=X^{k}-\eta_{k}\mathrm{msgn}(G^{k})$. Furthermore, convergence guarantees for Muon have been extensively studied (e.g., see (An et al., 2025; Chen et al., 2025; Kovalev, 2025; Li and Hong, 2025; Riabinin et al., 2025; Sato et al., 2025; Sfyraki and Wang, 2025; Shen et al., 2025)), and numerous new variants—such as SWAN (Ma et al., 2024), Scion (Pethick et al., 2025), Gluon (Riabinin et al., 2025), PolarGrad (Lau et al., 2025), Dion (Ahn and Xu, 2025; Ahn et al., 2025), and AdaMuon (Si et al., 2025)—have been proposed.
Beyond applying orthogonalization, a key innovation of Muon is its use of a GPU-friendly method—Newton-Schulz iterations (typically with five steps)—to perform inexact matrix orthogonalization, making it well-suited for modern foundation model training. In fact, several earlier methods, including spectral gradient descent (Carlson et al., 2015a; c; b) and orthogonalized gradient descent (Tuddenham et al., 2022), have already adopted SVD-based orthogonalization for matrix optimization problems. However, since SVD is not computationally efficient on GPUs when training large-scale neural networks, these methods fail to scale to foundation model training.
Inspired by Muon and its orthogonalization subroutine, we aim to develop a faster and more lightweight orthogonalization method to further enhance Muon and its variants. Specifically, our design leverages the widely observed phenomenon that the gradient matrices of NN parameters are often low-rank (see, e.g., (Hao et al., 2024; Hu et al., 2022; Malladi et al., 2023; Zhao et al., 2024)). To exploit this low-rank property, we propose performing low-rank orthogonalization by incorporating well-known low-rank matrix approximation techniques (Drineas et al., 2006a; Halko et al., 2011). Our approach first constructs a low-rank projection of the gradient matrix using QR decomposition on a sketched matrix, and then performs orthogonalization on the projected matrix by leveraging its structure. Our proposed low-rank orthogonalization offers two main advantages over existing orthogonalization methods:
  1. 1.
    Computational Efficiency: Traditional orthogonalization can be seen as computing the polar factor of a given full matrix. To exploit the low-rank property, our low-rank orthogonalization first computes the $Q$ factor of a smaller sketched matrix, and then computes the polar factor of a projected matrix constructed using the $Q$ factor. Since both the QR decomposition and polar decomposition are performed on much smaller matrices, our low-rank approach enjoys substantial computational savings for large-scale problems.
  2. 2.
    Noise Robustness: In the presence of noise, singular vectors associated with small singular values often vary significantly, leading to instability when directly applying orthogonalization to the full matrix. To circumvent this instability, our low-rank orthogonalization method clips these vectors to eliminate unreliable estimates, thereby stabilizing the orthogonalization process and yielding a robust estimate of the matrix sign.
These advantages will be illustrated in detail in Sections 3.1 and 4. Based on low-rank orthogonalization, we develop low-rank matrix-signed gradient descent (MSGD) and a low-rank variant of Muon. We also establish their complexity guarantees under mild assumptions.
Our main contributions are highlighted below.
  • We propose low-rank orthogonalization that can be readily incorporated into matrix-variate optimization algorithms such as Muon. It can be efficiently executed on GPUs and serves as a lightweight substitute for traditional orthogonalization methods that are directly applied to the full matrix, such as Newton-Schulz iterations.
  • Under mild assumptions, we establish the iteration complexity of low-rank MSGD and a low-rank variant of Muon. To the best of our knowledge, our results provide the first complexity guarantees for a broad class of inexact Muon-type algorithms, including vanilla Muon, under heavy-tailed noise.
The remainder of this paper is organized as follows. In Section 2, we introduce the notation and assumptions used throughout the paper. In Section 3, we propose low-rank orthogonalization and, based on it, develop low-rank MSGD and low-rank Muon. In Section 4, we present numerical results. Finally, we provide the proofs of the main results in Section 5.
2 Notation and Assumptions
[…]
We now make the following assumption throughout this paper.
Assumption 1.
  1. (a)
    There exists a finite $f_{\mathrm{low}}$ such that $f(X)\geq f_{\mathrm{low}}$ for all $X\in\mathbb{R}^{m\times n}$.
  2. (b)
    There exists an $L_{*}>0$ such that $\|\nabla f(X)-\nabla f(Y)\|_{*}\leq L_{*}\|X-Y\|$ for all $X,Y\in\mathbb{R}^{m\times n}$.
Assumption 1(a) is standard. Assumption 1(b) is natural in the analysis of Muon-type algorithms (e.g., see (Chen et al., 2025; Riabinin et al., 2025; Shen et al., 2025)). It follows from Assumption 1(b) that
$f(Y)\leq f(X)+\langle\nabla f(X),Y-X\rangle+\frac{L_{*}}{2}\|Y-X\|^{2}\qquad\forall X,Y\in\mathbb{R}^{m\times n}.$ (4)
We next provide a definition for approximate stationary points of problem (1).
Definition 1.
For any $\epsilon\in(0,1)$, we say that $X\in\mathbb{R}^{m\times n}$ is an $\epsilon$-nuclear norm stationary point (NSP) of problem (1) if it satisfies $\|\nabla f(X)\|_{*}\leq\epsilon$, and that it is an $\epsilon$-stochastic nuclear norm stationary point (SNSP) of problem (1) if it satisfies $\mathbb{E}[\|\nabla f(X)\|_{*}]\leq\epsilon$.
3 Matrix Optimization with Low-Rank Orthogonalization
In this section, we propose matrix optimization algorithms with low-rank orthogonalization for solving (1). In particular, we first propose low-rank orthogonalization in Section 3.1. Then, we propose low-rank MSGD in Section 3.2, and a low-rank variant of Muon in Section 3.3.
3.1 Low-Rank Orthogonalization
Orthogonalization techniques have attracted increasing attention in recent optimizer designs, as it has shown strong empirical performance in foundation model training (e.g., see (Bernstein and Newhouse, 2024; Jordan et al., ; Lau et al., 2025; Tuddenham et al., 2022)). In this subsection, we develop a low-rank orthogonalization method, leveraging low-rank matrix approximation techniques, that serves as a lightweight substitute for the existing orthogonalization subroutines used in matrix-variate optimizers.
Algorithm 1 A Low-Rank Orthogonalization Method
Input: matrix ${M}\in\mathbb{R}^{m\times n}$, rank trial $r\in\mathbb{Z}_{+}\cap[1,\varrho]$.
Output: approximate matrix sign $M_{O}\in\mathbb{R}^{m\times n}$.
Draw a Gaussian random matrix $G\in\mathbb{R}^{n\times r}$.
Perform a QR decomposition on $MG$ to obtain a column-orthogonal Q factor $Q\in\mathbb{R}^{m\times r}$.
Return $M_{O}=Q\mathrm{msgn}(Q^{T}M)$. (On GPUs, $\mathrm{msgn}(Q^{T}M)$ is recommended to be estimated via Newton-Schulz iterations.)
Specifically, our low-rank orthogonalization method, presented in Algorithm 1, is based on Gaussian sketching (Halko et al., 2011). This method first draws a Gaussian random matrix $G\in\mathbb{R}^{n\times r}$ with $r\ll\varrho$, and performs a QR decomposition on $MG$ to obtain a column-orthogonal Q factor $Q\in\mathbb{R}^{m\times r}$. Then, it computes $\mathrm{msgn}(Q^{T}M)\in\mathbb{R}^{r\times n}$ and returns $M_{O}=Q\mathrm{msgn}(Q^{T}M)$ as a low-rank approximation for $\mathrm{msgn}(M)$. As will be shown in Theorem 1, $M_{O}$ represents the matrix sign of $QQ^{T}M$, which is a low-rank approximation of $M$. Its proof is deferred to Section 5.1.
Theorem 1.
Consider Algorithm 1 with inputs $M\in\mathbb{R}^{m\times n}$ and $r\in\mathbb{Z}_{+}\cap[1,\varrho]$, where $\varrho:=\min\{m,n\}$. Let $Q\in\mathbb{R}^{m\times r}$ be generated by Algorithm 1. Then, for any $r_{*}$ satisfying $2\leq r_{*}\leq r-2$, it holds that
$\mathbb{E}[\|(I-QQ^{T})M\|_{F}]\leq\Big(1+\frac{r_{*}}{r-r_{*}-1}\Big)^{1/2}\|M-[M]_{r_{*}}\|_{F}.$ (5)
Moreover, we have $\mathrm{msgn}(QQ^{T}M)=Q\mathrm{msgn}(Q^{T}M)$.
Remark 1.
The relation (5) is adapted from (Halko et al., 2011, Theorem 10.5), where additional guarantees—such as those involving different matrix norms and high-probability bounds—can also be found. In addition, low-rank matrix approximation based on column selection (e.g., see (Drineas et al., 2006a; b)) can be used to develop a low-rank orthogonalization method. However, since its approximation guarantee is more complicated than that of the Gaussian sketching-based approach, we defer the column-selection-based method to the supplementary materials.
Next, we illustrate two major advantages of our low-rank orthogonalization method, namely, computational efficiency and noise robustness, through synthetic experiments on randomly generated matrices.
Computational Efficiency
We compare the computation time on GPUs for calculating inexact matrix sign of high-dimensional matrices using Newton-Schulz iterations, low-rank orthogonalization based on Gaussian sketching (Algorithm 1) and column selection (see supplementary materials), and truncated SVD.
For each $n\in\{1000,2000,5000,10000\}$, we generate 50 random matrices $M\in\mathbb{R}^{n\times n}$, with each entry drawn from the standard Gaussian distribution. We then apply all competing orthogonalization methods to estimate $\mathrm{msgn}(M)$. We implement Newton-Schulz iterations as provided in Muon (Jordan et al., ) with 5 iterations. For our low-rank orthogonalization methods, we set $r=0.1n$ as the input rank parameter, use command torch.linalg.qr to compute the Q factor, and apply 5 iterations of Newton-Schulz scheme following Muon (Jordan et al., ) to compute the matrix sign of $Q^{T}M$. In addition, we use command torch.pca_lowrank to efficient perform truncated SVD, setting the rank to $0.1n$ for truncation.
Figure 1: Left: Comparison of GPU computation time across Newton-Schulz iterations (NS), our low-rank orthogonalization with Gaussian sketching (GS) and column selection (CS), and truncated SVD (TSVD). Right: Speedup ratios of our low-rank orthogonalization methods compared to Newton-Schulz iterations.
Figure 2: Time distribution for low-rank orthogonalization with Gaussian sketching, including the QR decomposition, the Newton-Schulz iterations for computing the matrix sign, and other computations.
We present a comparison of the computation time on GPUs for all competing methods in Figure 1, and the time distribution of our low-rank orthogonalization method with Gaussian sketching in Figure 2. From Figure 1, we observe that our low-rank orthogonalization method significantly reduces computation time compared to Newton-Schulz iterations and truncated SVD. Although both truncated SVD and our low-rank orthogonalization exploit low-rank structure for computation, truncated SVD is not well-suited for GPU environments and is therefore slower than the Newton-Schulz iterations. From Figure 2, we observe that in our low-rank orthogonalization method with Gaussian sketching, the QR decomposition accounts for approximately half of the total time, while the Newton-Schulz iterations and other operations (such as generating Gaussian random matrices and performing matrix multiplications) each make up roughly half of the remaining time.
Noise Robustness
In addition to reducing computation time, our low-rank orthogonalization methods also produce more robust estimates of the matrix sign for low-rank matrices in the presence of noise. This robustness is particularly important in foundation model training, where gradients often exhibit low-rank structure. We now compare the performance of Newton–Schulz iterations and our low-rank methods in estimating the matrix sign of noisy, low-rank matrices.
For each $n\in\{1000,2000,5000,10000\}$, we first randomly generate 10 nearly low-rank matrices $M\in\mathbb{R}^{n\times n}$ following strategy: the top $0.1n$ singular values are set to $1$, the remaining singular values are set to $10^{-4}$, and the singular vectors are randomly generated orthogonal vectors. For each $M$, we generate 50 noise matrices $N\in\mathbb{R}^{n\times n}$, with each entry drawn from a Gaussian distribution with mean zero and variance $\sigma^{2}\in\{0.1,1,10\}$, and construct noisy matrices $M_{N}=M+N$. We next apply Newton-Schulz iterations, and our low-rank orthogonalization method based on Gaussian sketching (Algorithm 1), using an input rank parameter $0.1n$, to estimate $\mathrm{msgn}(M_{N})$. For each $M$, we compute the trace of the empirical covariance matrix of the estimates of $\mathrm{msgn}(M_{N})$ produced by both Newton-Schulz iterations and Algorithm 1. Both methods are implemented in the same way as in the above synthetic experiments to illustrate the computational efficiency.
Figure 3: Comparison of variance levels in matrix sign estimation across Newton-Schulz iterations (NS), and our low-rank orthogonalization methods using Gaussian sketching (GS), applied to nearly low-rank matrices in the presence of noise.
Figure 3 presents a comparison of the average trace of the covariance matrices across all competing methods for matrix sign estimation. From this figure, we observe that, compared to applying the Newton-Schulz iterations to the full matrix, our low-rank orthogonalization method yields matrix sign estimates that are significantly less sensitive to noise. This is because singular vectors associated with small singular values are more sensitive to noise than others, so removing them improves the robustness of the matrix sign computation. As a result, our low-rank method provides more stable and robust estimates of the matrix sign function, particularly for nearly low-rank matrices.
3.2 Low-Rank Matrix-Signed Gradient Descent
In this subsection, we propose low-rank MSGD methods, including a fixed-rank variant and a safeguarded variant with adaptive ranks. Both variants use low-rank orthogonalization as a subroutine.
Algorithm 2 Low-Rank MSGD
Input: starting point $X^{0}\in\mathbb{R}^{m\times n}$, rank parameter $r\in\mathbb{Z}_{+}\cap[1,\varrho]$, step sizes $\{\eta_{k}\}\subset(0,\infty)$.
for $k=0,1,2,\ldots$ do
  Call Algorithm 1 (see Section 3.1) with $(M,r)=(\nabla f(X^{k}),r)$ to obtain an approximate matrix sign $M^{k}_{O}$, such that $M^{k}_{O}=\mathrm{msgn}(M^{k}_{Q})$, where $M^{k}_{Q}$ is a low-rank approximation for $\nabla f(X^{k})$.
  Update the next iterate: $X^{k+1}=X^{k}-\eta_{k}M^{k}_{O}$.
end for
At each iteration $k\geq 0$, the fixed-rank variant invokes Algorithm 1 to compute $M^{k}_{O}=\mathrm{msgn}(M_{Q}^{k})$, where $M_{Q}^{k}$ is a low-rank approximation of $\nabla f(X^{k})$. Then, the next iterate $X^{k+1}$ is obtained by performing a line search update from $X^{k}$ along the matrix-signed direction $M^{k}_{O}$. Details of this method are presented in Algorithm 2.
The following theorem provides a convergence guarantee for Algorithm 2, whose proof is deferred to Section 5.2.
Theorem 2.
Suppose that Assumption 1 holds. Let $\{(X^{k},M_{O}^{k})\}$ be the sequence generated by Algorithm 2 with $M_{O}^{k}=\mathrm{msgn}(M_{Q}^{k})$ for all $k\geq 0$ and step sizes $\{\eta_{k}\}$ given by $\eta_{k}=(k+1)^{-1/2}$ for all $k\geq 0$. Then, it holds that for all $K\geq 3$,
$\min_{0\leq k\leq K-1}\{\|\nabla f(X^{k})\|\}\leq\frac{f(X^{0})-f_{\mathrm{low}}+L_{*}\ln K}{K^{1/2}}+\frac{2}{K^{1/2}}\sum_{k=0}^{K-1}\frac{\|\nabla f(X^{k})-M_{Q}^{k}\|_{*}}{(k+1)^{1/2}},$ (6)
where $L_{*}$ is defined in Assumption 1.
Remark 2.
From Theorem 2, we observe that when $\{(X^{k},M_{Q}^{k})\}$ satisfies $\sum_{k=0}^{K-1}\|\nabla f(X^{k})-M_{Q}^{k}\|_{*}(k+1)^{-1/2}=\widetilde{\mathcal{O}}(1)$, it holds that $\min_{0\leq k\leq K-1}\{\|\nabla f(X^{k})\|_{*}\}=\widetilde{\mathcal{O}}(K^{-1/2})$, which matches, up to a logarithmic factor, the well-established optimal convergence rate for nonconvex optimization (e.g., see (Carmon et al., 2020)). Moreover, if the gradients $\{\nabla f(X^{k})\}$ have low rank when $k$ is sufficiently large (as is often the case during deep neural network training (Zhao et al., 2024)), we can tune the rank parameter $r$ to be close to the effective rank of the gradients, causing the sequence $\{\|\nabla f(X^{k})-M_{Q}^{k}\|_{*}\}$ to remain close to zero.
We next describe a safeguarded low-rank MSGD method with adaptively updated ranks. At each iteration $k\geq 0$, this method invokes Algorithm 1 with $(M,r)=(\nabla f(X^{k}),r_{k})$ to obtain $M^{k}_{O}=\mathrm{msgn}(M_{Q}^{k})$, where $M_{Q}^{k}$ is a low-rank approximation of $\nabla f(X^{k})$ such that the approximation error satisfies (7). Then, the next iterate $X^{k+1}$ is obtained by performing a line search update from $X^{k}$ along the matrix-signed direction $M^{k}_{O}$, with a suitable step size. Details of this method are provided in Algorithm 3.
It is noteworthy that the approximation error in (7) holds for some $r_{k}\leq\varrho$ because when $r_{k}=\varrho$, the error $\|M^{k}_{Q}-\nabla f(X^{k})\|_{*}$ is zero. When the matrix is low-rank or has a small effective rank, $r_{k}$ can be chosen to be much smaller than $\varrho$. In practice, one can gradually increase the trial ranks to find $r_{k}$ such that (7) holds.
Algorithm 3 Safeguarded Low-Rank MSGD
Input: starting point $X^{0}\in\mathbb{R}^{m\times n}$, initial rank trial $r_{0}\in\mathbb{Z}_{+}\cap[1,\varrho]$, step sizes $\{\eta_{k}\}\subset(0,\infty)$, control errors $\{\delta_{k}\}\subset(0,\infty)$
for $k=0,1,2,\ldots$ do
  Call Algorithm 1 (see Section 3.1) with $(M,r)=(\nabla f(X^{k}),r_{k})$ to obtain an approximate matrix sign $M^{k}_{O}$ such that $M_{O}^{k}=\mathrm{msgn}(M_{Q}^{k})$ and
$\|\nabla f(X^{k})-M_{Q}^{k}\|_{*}\leq\delta_{k}.$ (7)
  Update the next iterate: $X^{k+1}=X^{k}-\eta_{k}M^{k}_{O}$.
end for
The following theorem establishes an iteration complexity of Algorithm 3, whose proof is deferred to Section 5.2.
Theorem 3.
Suppose that Assumption 1 holds. Let $f_{\mathrm{low}}$ and $L_{*}$ be given in Assumption 1, and define
$U_{\mathrm{gd}}:=f(X^{0})-f_{\mathrm{low}}+L_{*}+4.$ (8)
Let $\{X^{k}\}$ be generated by Algorithm 3 with inputs $\{(\eta_{k},\delta_{k})\}$ given by $\eta_{k}=\delta_{k}={(k+1)^{-1/2}}$ for all $k\geq 0$. Then, for any $\epsilon\in(0,1)$, it holds that $\min_{0\leq k\leq K-1}\{\|\nabla f(X^{k})\|_{*}\}\leq\epsilon$ for all $K$ satisfying
$K\geq\max\Big\{\Big(\frac{4U_{\mathrm{gd}}}{\epsilon}\ln\Big(\frac{4U_{\mathrm{gd}}}{\epsilon}\Big)\Big)^{2},3\Big\}.$
Remark 3.
From Theorem 3, we observe that Algorithm 3 achieves an iteration complexity of $\widetilde{\mathcal{O}}(\epsilon^{-2})$ for finding an $\epsilon$-NSP of (1). This complexity bound matches, up to a polylogarithmic factor, the lower complexity bound as established in (Carmon et al., 2020).
3.3 Low-Rank Muon
In this subsection, we propose a low-rank variant of Muon (Jordan et al., ), and analyze its iteration complexity under heavy-tailed noise.
This method follows a framework similar to Muon (Jordan et al., ), but instead of directly computing the matrix sign of the momentum update, it computes only its low-rank approximation. Specifically, at each iteration of this method, it first performs a momentum update to generate $M^{k}$ by aggregating stochastic gradients of $f$ evaluated at $X^{0},\ldots,X^{k}$. Then, Algorithm 1 is invoked to obtain $M^{k}_{O}=\mathrm{msgn}(M_{Q}^{k})$, where $M_{Q}^{k}$ is a low-rank approximation of $\nabla f(X^{k})$ such that the approximation error satisfies (10). The next iterate $X^{k+1}$ is obtained by performing a line search update from $X^{k}$ along the matrix-signed direction $M_{O}^{k}$ with a suitable step size. Details of this method are given in Algorithm 4.
Before analyzing the complexity of Algorithm 4 for computing an approximate solution to (1), we make the following heavy-tailed noise assumption regarding the stochastic gradient $G(\cdot;\xi)$.
Assumption 2.
The stochastic gradient estimator $G:\mathbb{R}^{m\times n}\times\Xi\to\mathbb{R}^{m\times n}$ satisfies
$\mathbb{E}[G(X;\xi)]=\nabla f(X),\quad\mathbb{E}[\|G(X;\xi)-\nabla f(X)\|_{F}^{\alpha}]\leq\sigma^{\alpha}$
for some $\sigma>0$ and $\alpha\in(1,2]$.
Algorithm 4 Low-Rank Muon
Input: starting point $X^{0}\in\mathbb{R}^{m\times n}$, initial rank trial $r_{0}\in\mathbb{Z}_{+}\cap[1,\varrho]$, step sizes $\{\eta_{k}\}\subset(0,\infty)$, weighting parameters $\{\theta_{k}\}\subset(0,1]$, control errors $\{\delta_{k}\}\subset(0,\infty)$.
Initialize: $M^{-1}=\mathbf{0}_{m\times n}$ and $\theta_{-1}=1$.
for $k=0,1,2,\ldots$ do
  Compute the full-rank search direction:
$M^{k}=(1-\theta_{k-1})M^{k-1}+\theta_{k-1}G(X^{k};\xi^{k}).$ (9)
  Call Algorithm 1 (see Section 3.1) with $(M,r)=(M^{k},r_{k})$ to obtain an approximate matrix sign $M^{k}_{O}$ such that $M_{O}^{k}=\mathrm{msgn}(M^{k}_{Q})$ and
$\|M^{k}-M_{Q}^{k}\|_{*}\leq\delta_{k}.$ (10)
  Update the next iterate:
$X^{k+1}=X^{k}-\eta_{k}M_{O}^{k}.$ (11)
end for
The following theorem establishes the iteration complexity of Algorithm 4, whose proof is deferred to Section 5.3.
Theorem 4.
Suppose that Assumptions 1 and 2 hold. Let $f_{\mathrm{low}}$ and $L_{*}$ be given in Assumption 1, and $\sigma$ and $\alpha$ be given in Assumption 2. Define
$U_{\mathrm{mn}}$ $:=f(X^{0})-f_{\mathrm{low}}+\sigma^{\alpha}+2L_{*}+4+2(\alpha-1)(2\varrho^{1/2}/\alpha)^{\alpha/(\alpha-1)}+6L_{*}^{\alpha}+4\sigma^{\alpha}.$ (12)
Let $\{X^{k}\}$ be generated by Algorithm 4 with input parameters $\{(\eta_{k},\theta_{k},\delta_{k})\}$ given by $\eta_{k}=(k+1)^{-(2\alpha-1)/(3\alpha-2)}$, $\theta_{k}=(k+1)^{-\alpha/(3\alpha-2)}$, and $\delta_{k}=(k+1)^{-(\alpha-1)/(3\alpha-2)}$ for all $k\geq 0$. Then, for any $\epsilon\in(0,1)$, it holds that $\mathbb{E}[\|\nabla f(X^{\iota_{K}})\|_{*}]\leq\epsilon$ for all $K$ satisfying
$K\geq\max\Big\{\Big(\frac{2(3\alpha-2)U_{\mathrm{mn}}}{(\alpha-1)\epsilon}\ln\Big(\frac{2(3\alpha-2)U_{\mathrm{mn}}}{(\alpha-1)\epsilon}\Big)\Big)^{\frac{3\alpha-2}{\alpha-1}},3\Big\},$
where $\iota_{K}$ is uniformly drawn from $\{0,\ldots,K-1\}$.
Remark 4.
From Theorem 4, one can observe that Algorithm 4 achieves an iteration complexity of $\widetilde{\mathcal{O}}(\epsilon^{-(3\alpha-2)/(\alpha-1)})$ for finding an $\epsilon$-SNSP of problem (1), which matches the optimal dependence on $\epsilon$ in the complexity results for vector-variate stochastic first-order methods under heavy-tailed noise (see, e.g., (He et al., 2025; Zhang et al., 2020)). To the best of our knowledge, our result is the first to show that the inexact Muon-type algorithms can achieve an iteration complexity with optimal dependence on $\epsilon$ when applied to matrix-variate optimization under heavy-tailed noise.
4 Numerical Experiments
In this section, we present numerical experiments to evaluate the performance of low-rank Muon (Algorithm 4) and compare our method with vanilla Muon, AdamW, and SGD. The experiments are conducted on GPT-2 pretraining (Section 4.1) and LLaMA pretraining (Section 4.2). All experiments are executed on a server with two NVIDIA A100 GPUs (80 GB). The code to implement the proposed algorithm on these numerical examples is available at https://github.com/dengzhanwang/Low-rank-Muon.
4.1 GPT-2 Pretraining
In this subsection, we consider pretraining GPT-2 (Radford et al., 2019), a transformer-based language model. We experiment with GPT-2 models of sizes 60M, 135M, 350M, and 1B on the same three datasets tested in the Muon GitHub repository (Jordan et al., ): FineWeb10B, FineWeb100B, and FineWebEdu10B.
Table 1: Comparison of validation perplexity and computational time for all competing methods on GPT-2 pretraining.
Dataset Method Validation Perplexity Computational Time
60M 135M 350M 1B 60M 135M 350M 1B
FineWeb10B SGDM 48.85 48.85 31.54 27.86 2.51e3 7.08e3 1.20e4 2.70e5
AdamW 43.56 43.56 25.72 21.58 2.52e3 7.08e3 1.21e4 2.69e5
GaLore 57.61 39.59 30.87 28.58 2.73e3 7.12e3 1.24e4 2.76e5
SOAP 41.32 55.37 38.45 34.58 2.56e3 7.09e3 1.23e4 2.73e5
Muon 32.89 32.99 28.48 20.99 2.76e3 7.57e3 1.32e4 3.04e5
LR-Muon100 35.21 28.83 27.32 21.70 2.69e3 7.49e3 1.25e4 2.90e5
LR-Muon200 33.98 27.34 25.64 20.69 2.79e3 7.54e3 1.26e4 2.91e5
FineWeb100B SGDM 51.48 51.48 77.50 29.75 2.69e3 7.25e3 1.04e4 2.52e5
AdamW 43.52 43.51 27.47 23.53 2.70e3 7.21e3 1.05e4 2.56e5
GaLore 56.42 58.09 31.19 26.53 2.81e3 7.31e3 1.15e4 2.67e5
SOAP 41.32 44.32 30.98 25.56 2.71e3 7.23e3 1.07e4 2.58e5
Muon 34.04 33.02 32.76 20.75 2.79e3 7.70e3 1.19e4 3.09e5
LR-Muon100 35.78 30.19 27.59 21.76 2.77e3 7.62e3 1.09e4 2.87e5
LR-Muon200 34.02 28.31 25.86 20.45 2.80e3 7.69e3 1.10e4 2.91e5
FineWebEdu10B SGDM 50.92 77.50 31.54 25.49 2.56e3 7.09e3 1.20e4 2.53e5
AdamW 43.56 26.45 21.01 22.45 2.59e3 7.09e3 1.20e4 2.56e5
GaLore 69.49 39.59 57.42 37.45 2.64e3 7.12e3 1.24e4 2.60e5
SOAP 55.37 41.77 47.90 39.20 2.60e3 7.11e3 1.22e4 2.58e5
Muon 32.88 23.89 22.23 19.54 2.80e3 7.63e3 1.33e4 3.05e5
LR-Muon100 35.60 23.50 22.39 19.97 2.74e3 7.53e3 1.26e4 2.85e5
LR-Muon200 33.93 22.37 20.96 18.85 2.83e3 7.59e3 1.29e4 2.87e5
We apply low-rank Muon, Muon, SOAP, GaLore, AdamW, and SGD with momentum (SGDM) for GPT-2 pretraining. We implement two low-rank Muon variants with rank parameters 100 and 200, abbreviated as LR-Muon100 and LR-Muon200, respectively. Following the experiments in (Jordan et al., 2024; Lau et al., 2025), we use AdamW to train the embedding and head layers for both Muon and low-rank Muon. We initialize all methods with the same weights from pretrained GPT-2 models, and terminate all methods after one training epoch, consisting of 5000 iterations. We compare all methods using validation perplexity, a standard metric in foundation model training (e.g., (Radford et al., 2019; Touvron et al., 2023; Vaswani et al., 2017)), which is defined as the exponential of the validation loss and serves to amplify performance differences. The algorithmic parameters are carefully tuned to suit each method well in terms of computational performance. More details about the experimental setups, including algorithm and GPT-2 model parameters, are provided in supplementary materials.
Figure 4: Comparison of validation perplexity versus computational time for all competing methods in training GPT-2 models.
We present a comparison of validation perplexity and computational time in Table 1. From this table, we observe that for GPT-2 with a model size of 60M, our low-rank Muon achieves slightly worse validation perplexity than full-rank Muon, but performs better than AdamW, SGDM, SOAP, and GaLore. For GPT-2 models with larger sizes, our low-rank Muon achieves better validation perplexity than all other competing methods. We also observe that the computational time of our low-rank Muon improves upon that of vanilla Muon, although all Muon-type methods remain slower than SGDM and AdamW. These observations show that our low-rank orthogonalization enables obtaining more generalizable GPT-2 models when the model size is large, but may be less effective for smaller models. In addition, our low-rank orthogonalization saves computational time compared to orthogonalization by Newton-Schulz iterations, but the gain is not as significant as those presented in Section 3.1, since the forward and backward passes of neural networks account for the majority of the computational time.
We plot the convergence behavior of validation perplexity versus training iterations in Figure 4. From this figure, we observe that for each model size, the competing methods exhibit similar performance patterns across the three datasets. For the model size of 60M, all Muon-type methods consistently outperform AdamW and SGDM, while our low-rank Muon performs slightly worse than vanilla Muon. For the 135M model, our low-rank Muon outperforms vanilla Muon and other training methods. Vanilla Muon performs similarly to AdamW, and all methods show a large improvement compared to SGDM. For the model size of 350M, our low-rank Muon has slightly better performance than AdamW, while outperforming Muon and SGDM by a larger margin. These observations support that our low-rank Muon improves upon vanilla Muon in GPT-2 training and is more effective for models with larger sizes.
4.2 LLaMA Pretraining
In this subsection, we consider pretraining LLaMA (Touvron et al., 2023), a transformer-based language model with a more refined architecture than GPT-2. We experiment with LLaMA models of sizes 60M, 135M, 350M, and 1B on the same three datasets tested in the Muon GitHub repository (Jordan et al., ): FineWeb10B, FineWeb100B, and FineWebEdu10B.
Table 2: Comparison of validation perplexity and computational time for all competing methods on LLaMA pretraining.
Dataset Method Validation Perplexity Computational Time
60M 135M 350M 1B 60M 135M 350M 1B
FineWeb10B SGDM 94.90 86.99 86.49 44.71 $2.49\text{e3}$ $5.27\text{e3}$ $1.20\text{e4}$ $2.77\text{e5}$
AdamW 50.51 40.17 35.76 24.14 $2.46\text{e3}$ $5.27\text{e3}$ $1.21\text{e4}$ $2.78\text{e5}$
GaLore 94.90 86.99 86.49 44.71 $2.49\text{e3}$ $5.27\text{e3}$ $1.20\text{e4}$ $2.77\text{e5}$
SOAP 50.51 40.17 35.76 24.14 $2.46\text{e3}$ $5.27\text{e3}$ $1.21\text{e4}$ $2.78\text{e5}$
Muon 36.68 37.96 36.43 20.99 $2.59\text{e3}$ $5.69\text{e3}$ $1.35\text{e4}$ $3.18\text{e5}$
LR-Muon100 37.33 34.08 33.54 21.70 $2.69\text{e3}$ $5.66\text{e3}$ $1.28\text{e4}$ $2.79\text{e5}$
LR-Muon200 36.05 34.37 32.37 20.69 $2.70\text{e3}$ $5.70\text{e3}$ $1.30\text{e4}$ $2.91\text{e5}$
FineWeb100B SGDM 81.59 77.50 87.49 37.56 $2.60\text{e3}$ $5.44\text{e3}$ $1.05\text{e4}$ $2.65\text{e5}$
AdamW 41.49 40.54 35.47 23.89 $2.61\text{e3}$ $5.47\text{e3}$ $1.06\text{e4}$ $2.69\text{e5}$
GaLore 58.97 45.14 37.05 35.56 $2.65\text{e3}$ $5.50\text{e3}$ $1.09\text{e4}$ $2.69\text{e5}$
SOAP 55.08 52.28 47.52 43.89 $2.64\text{e3}$ $5.50\text{e3}$ $1.07\text{e4}$ $2.70\text{e5}$
Muon 38.49 37.66 36.66 21.76 $2.58\text{e3}$ $5.63\text{e3}$ $1.07\text{e4}$ $3.16\text{e5}$
LR-Muon100 37.68 34.58 33.71 22.86 $2.63\text{e3}$ $5.55\text{e3}$ $1.10\text{e4}$ $2.79\text{e5}$
LR-Muon200 36.26 34.20 32.73 20.98 $2.75\text{e3}$ $5.69\text{e3}$ $1.11\text{e4}$ $2.88\text{e5}$
FineWebEdu10B SGDM 82.63 73.48 73.48 47.56 $2.57\text{e3}$ $5.27\text{e3}$ $1.20\text{e4}$ $2.76\text{e5}$
AdamW 41.82 32.96 29.12 22.54 $2.60\text{e3}$ $5.29\text{e3}$ $1.20\text{e4}$ $2.78\text{e5}$
GaLore 64.04 68.13 45.13 43.56 $2.57\text{e3}$ $5.27\text{e3}$ $1.20\text{e4}$ $2.76\text{e5}$
SOAP 40.32 28.10 27.79 24.54 $2.60\text{e3}$ $5.29\text{e3}$ $1.20\text{e4}$ $2.78\text{e5}$
Muon 28.57 30.72 29.47 22.67 $2.68\text{e3}$ $5.84\text{e3}$ $1.33\text{e4}$ $3.25\text{e5}$
LR-Muon100 30.52 27.72 27.24 21.36 $2.72\text{e3}$ $5.66\text{e3}$ $1.24\text{e4}$ $2.88\text{e5}$
LR-Muon200 29.29 27.65 26.87 19.54 $2.81\text{e3}$ $5.70\text{e3}$ $1.26\text{e4}$ $2.98\text{e5}$
Figure 5: Comparison of validation perplexity and computational time for all competing methods in training LLaMA models.
We apply low-rank Muon, Muon, AdamW, GaLore, SOAP and SGD with momentum (SGDM) for pretraining LLaMA. Similar to Section 4.1, we implement two versions of low-rank Muon with rank parameters $100$ and $200$, and we use AdamW to train the embedding and head layers for both Muon and low-rank Muon. We initialize all methods with the same weights from pretrained LLaMA models, and terminate all methods after one training epoch, consisting of 5000 iterations. We compare all methods using validation perplexity. The algorithmic parameters are carefully tuned to suit each method well in terms of computational performance. More details about the experimental setups, including algorithm and LLaMA model parameters, are provided in supplementary materials.
We present a comparison of validation perplexity and computational time in Table 2. From this table, we observe that our low-rank Muon consistently achieves better validation perplexity compared to Muon, GaLore, and SOAP for the majority of model sizes and tested datasets, while vastly improving upon the validation perplexity achieved by AdamW and SGDM. We also observe that the computational time of our low-rank Muon improves upon that of vanilla Muon, though all Muon-type methods remain slower than SGDM and AdamW. These observations demonstrate that our low-rank orthogonalization produces more generalizable LLaMA models across all tested model sizes and datasets. Moreover, it reduces computational time compared to the Newton-Schulz iterations for orthogonalization, although the improvement is less significant than the gains discussed in Section 3.1, as most of the computational time is dominated by the forward and backward passes of the neural networks.
We plot the convergence behavior of validation perplexity versus training iterations in Figure 5. From this figure, we observe that for the model size of 60M, all Muon-type methods significantly outperforms AdamW and SGDM, while our low-rank Muon has similar performance compared to the vanilla Muon. For the 135M and 350M model size, our low-rank Muon significantly outperforms the vanilla Muon and other training methods. These observations support that our low-rank Muon consistently outperforms all other methods for training LLaMA models.